p=2c+d/3
Is this the correct answer below solving for d?
d= -6c+3p

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

p=2c+d/3
Is this the correct answer below solving for d?
d= -6c+3p
-------------------
If your problem is really
p = (2c+d)/3
---
3p = 2c+d
---
Then d = 3p-2c
=========================
If your problem is really
p = 2c + (d/3)
3p = 6c + d
Then d = 3p-6p
=========================
Cheers,
Stan H.
=============

Answer by rfer(16322)

p=2c+d/3
subtract 2c from both sides
p-2c=d/3
multiply both sides by 3
3p-6c=d
your work looks good.